OpenStudy (anonymous):

Use properties of real numbers to rewrite the expresion. (3/5) (5/3) 8/15 15/8 0 1

OpenStudy (anonymous):

thank you! :)

OpenStudy (anonymous):

It would be 1 since the there both reciprocals of each other.

OpenStudy (anonymous):

3/5 X 5/3 = 15/15 = 1